OpenStudy (anonymous):

What is the solution of the system of equations? -3x-4y-3z=-7 2x-6y+2z=3 5x-2y+5z=9 Please help me I'm really confused!

OpenStudy (amoodarya):

multiply by (-) the 2nd eq then plus 1st eq so you can find y

OpenStudy (anonymous):

what? i'm still confused, multiply what in the second equation by (-)? @amoodarya

OpenStudy (amoodarya):

hint the system of equations has no answer

OpenStudy (amoodarya):

-3x-4y-3z=-7 -2x+6y-2z=-3 *(-1) 5x-2y+5z=9

OpenStudy (amoodarya):

sum of 1st and 2nd -3x-4y-3z=-7 -2x+6y-2z=-3 is -5x+2y-5z=-10

OpenStudy (amoodarya):

now compare with 3rd eq

OpenStudy (anonymous):

wait so there is no solution? then why do you have to compare them? @amoodarya

OpenStudy (amoodarya):

-5x+2y-5z=-10 5x-2y+5z=9 it is impossible try to compare them

OpenStudy (amoodarya):

sum of them is 5x-2y+5z=9 -5x+2y-5z=-10 0x+0y+0z=9-10 so 0=-1 the system has no answer

OpenStudy (anonymous):

oh okay I get it. thanks! can you help me with another one?

OpenStudy (anonymous):

@amoodarya

OpenStudy (amoodarya):

yes

OpenStudy (anonymous):

Solve the system by elimination. -2x+2y+3z=0 -2x-y+z=-3 2x+3y+3z=5 I need help like step by step because I don't understand how to do this one at all @amoodarya

OpenStudy (amoodarya):

first name them (1) -2x+2y+3z=0 (2) -2x-y+z=-3 (3) 2x+3y+3z=5

OpenStudy (amoodarya):

no calculate 1+3 and 2+3

OpenStudy (anonymous):

so the first equation plus the second equation and the second equation plus the third equation?

OpenStudy (amoodarya):

now calculate eq 1+eq 3 and eq 2+eq 3 because "x" omited

OpenStudy (amoodarya):

you must have 5y+6z=5 4y+4z=2 ok ?

OpenStudy (anonymous):

Okay @amoodarya

OpenStudy (amoodarya):

sorry 5y+6z=5 2y+4z=2 now multiply the 1st by -4 multiply the 2nd by +6 then sum of them

OpenStudy (amoodarya):

then you must have -8y=-8 ok ?

OpenStudy (anonymous):

right ok @amoodarya

OpenStudy (anonymous):

wait how did you get -8y=-8?

OpenStudy (anonymous):

maybe @hero can help?

OpenStudy (amoodarya):

and now what is problem?

OpenStudy (amoodarya):

5y+6z=5 2y+4z=2 -4 *(5y+6z=5) +6* (2y+4z=2) -20y-24z=-20 12y+24z=12 sum of them -8y=-8 so y=1 canyou go on ?

OpenStudy (amoodarya):

then put y=1 you must have z=0 and then put y=1 and z=0 you must have x=1

OpenStudy (amoodarya):

Briana_Brooke are u there ?

OpenStudy (anonymous):

oh okay I see how you got it now is that it or is there more to do? @amoodarya

OpenStudy (amoodarya):

nothing more

OpenStudy (anonymous):

okay thanks!

hero (hero):

@Briana_Brooke

hero (hero):

This is the original system: -3x-4y-3z=-7 2x-6y+2z=3 5x-2y+5z=9

hero (hero):

Notice that the first and third terms in each equation are the same.

hero (hero):

Oh...I see...nevermind

hero (hero):

I didn't realize you guys went over two problems.

hero (hero):

Good job @amoodarya